Circular No. 6629

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 P. A. Caraveo, Istituto di Fisica Cosmica (IFC), CNR, Milan; R.

Mignani, Max-Planck-Institut für Extraterrestrische Physik, Garching; M. Tavani, Columbia University and IFC; and G. F. Bignami, Agenzia Spaziale Italiana and IFC, communicate: “Analysis of the Hubble Space Telescope (HST) PC set of images of GRB 970228 shows that the faint point-like object reported by Sahu et al. (IAUC 6619, 6606) has significant proper motion. By comparing the Mar. 26.2 and Apr. 7.2 images, we obtain the following relative displacements: Δα = +0.29 ± 0.09 pixels, Δδ = −0.26 ± 0.09 pixels, for a total displacement of 18 mas to the southeast, corresponding to a proper motion of 540 mas/yr. Five field stars all show total displacements < 0.09 pixels. Therefore, the object appears to be galactic. For example, it could be a compact star at 100 pc moving at 260 km/s. Its association with the reported optical transient and GRB 970228 is compelling. Further HST observations are urgently requested.”

SUPERNOVA 1997bq IN NGC 3147 E. Waagen, AAVSO, reports the following photometry by R. Zissell

(South Hadley, MA) for SN 1997bq on Apr. 15.237 UT: B = 14.68 ± 0.09, V = 14.65 ± 0.05, R = 14.58 ± 0.04, I = 14.58 ± 0.04. Zissell made an independent discovery of this object.

Visual magnitude estimate by S. Pesci, Milan, Italy: Apr. 12.8 UT, 14.8.

SUPERNOVA 1997bl IN ANONYMOUS GALAXY D. Hardin, CEA, Saclay, reports that the spectrum reported on IAUC

6605 has been fully reduced by S. Benetti, indicating that SN 1997bl is a type-Ia supernova with z = 0.19.

COMET C/1995 O1 (HALE-BOPP) Naked-eye total magnitude estimates: Apr. 3.04 UT, −0.5 (J. Bortle,

Stormville, NY); 4.78, −1.1 (K. Sárneczky, Ráktanya, Hungary); 6.90, −0.6 (B. H. Granslo, Tryvann, Oslo, Norway); 9.20, −0.5 (C. E. Spratt, Victoria, BC); 10.80, −0.8 (J. M. Trigo, Benicassim, Castellon, Spain); 11.85, −0.8 (A. Pereira, Cabo da Roca, Portugal); 14.82, −0.5 (M. V. Zanotta, Milan, Italy); 16.85, −0.4 (R. J. Bouma, Groningen, The Netherlands).

1997 April 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green